Which number is closest to the slope of the line: 3y+2x-13=0

A. -2/3
B. 1
C. 1/4
D. 3/2

Answer by tutorcecilia(2152)

3y+2x-13=0 [re-write in the slope-intercept (y=mx+b) format by solving for the y-term]
.
3y+2x-2x-13=0-2x
3y-13=-2x
3y-13+13=-2x+13
3y=-2x+13
3y/3=-2x/3+13/3
y=-2/3x+13/3
Slope=m = -2/3